OpenStudy (anonymous):

write and solve an equation to find value of variable 104,137,154,131, x; mean = 130

OpenStudy (amistre64):

how would you go about finding the mean of a set of data like this to begin with? spose that x is already some known value

OpenStudy (anonymous):

add i believe

OpenStudy (amistre64):

adding is a good step, after you add them alltogether, then you would want to divide by something right? what would we want to divide it by?

OpenStudy (anonymous):

4 or 5

OpenStudy (amistre64):

hmm, lets stick to 5 since we want to count x as a known value :) so, this gives us \[\frac{104+137+154+131+x}{5} = 130\] now, how would we go about getting x all by itself?

OpenStudy (anonymous):

so whatever i get i just put x as 1

OpenStudy (anonymous):

OMG OMG OMG, Im on teem Jake also! *virtual high five*

OpenStudy (anonymous):

haha

OpenStudy (amistre64):

no, x is unknown so it makes no sense to just put x as 1; but since we have an idea what to do with x if we DID know it; then this is the setup that we would take

OpenStudy (amistre64):

once we know the setup, it should be pretty basic to rearrange the equation to solve for x

OpenStudy (anonymous):

ok could you help put it in the order

OpenStudy (amistre64):

sure, how do we undo that 5 in the denominator?

OpenStudy (anonymous):

well since we divided didnt we use it already

OpenStudy (anonymous):

i dont know if its right but i got 127

OpenStudy (amistre64):

we havent done any math tey, we have just taken the information and put it into a format that we can work with as opposed to some abstract wordings. \[\frac{104+137+154+131+x}{5} = 130\] to get to the x, lets multiply everything by 5 \[\cancel{5}*\frac{104+137+154+131+x}{\cancel{5}} = 130*5\] \[104+137+154+131+x = 130*5\] now its just a matter of subtracting the knowns from the left and right in order to evaluate x

OpenStudy (anonymous):

so we subtract now

OpenStudy (amistre64):

correct

OpenStudy (anonymous):

just the basic numbers together?

OpenStudy (amistre64):

yes; if you want to add up all the numbers and then subtract the total amount from each side, thats fine

OpenStudy (amistre64):

what is: 104+137+154+131 ?

OpenStudy (anonymous):

526

OpenStudy (amistre64):

then x should be 130*5 - 526

OpenStudy (anonymous):

124 k great i get it now thanks alot

OpenStudy (amistre64):

\[104+137+154+131+x = 130*5\] \[526+x = 130*5\] \[x = 130*5-526\]

OpenStudy (amistre64):

good luck